# Flag complex face structures and decompositions

### v1

@@ 2.2 Colorings, Tchebyshev subdivisions, and intervals
$[\alpha^{k}]P_{\Delta}(2\alpha+1)$ $=\sum_{j=k}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)2^{k}\binom{j}{k}$
$=2^{k}\sum_{j=k}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)\binom{j}{k}.$
We claim that the $f$-vector of the complex $\mathcal{D}(\Gamma)$ from Definition 2.19 has dimensions given by the sum above. Given a $\left(\frac{d}{2}-j-1\right)$-dimensional $F\in\Gamma$, there are $j$ unused colors forming the set $C_{F}\subset\left[\frac{d}{2}\right]$. Choosing $k$ of them gives the subset $Q\subset C_{F}$ from Definition 2.19 and is counted by the term $f_{\frac{d}{2}-j-1}(\Gamma)\binom{j}{k}$ in the sum. Assigning a $\pm$ signs to each element of $Q$ gives the multiplication by $2^{k}$.
@@
(Definition 3.2 on p. 578) Given a cell complex (Definition 2.1) $A$ of dimension $D-1$, the $F$-polynomial of $A$ is
@@
Given a cell complex $A$ such that $(k-1)$-dimensional faces have $k$ vertices, we have that
@@
for a cell complex $T$ with $n$ vertices such that $(k-1)$-dimensional faces have $k$ vertices.
@@
(Cell complex/multicomplex generalization of Definition 2.1 on p. 575 of [18])
Let $A$ be a cell complex from Definition 2.1 or the (dual of) a multicomplex from Definition 2.2.
Given a cell complex $A$ on a vertex set $V$, fix a linear order $<$ on $V$ and introduce a new element $\widehat{1}$ larger than all elements of $V$. The Tchebyshev triangulation $T_{<}A$ of $A$ is the cell complex on the vertex set
@@
whose $(k-1)$-faces are all sets $\{(u_{1},v_{1}),\ldots,(u_{k},v_{k})\}$ satisfying $v_{1}<\cdots<v_{k}$ and the following conditions:
@@
Since the $f$-vector of $T_{<}(A)$ is independent of the ordering $<$, we will often drop the ordering and simply write $T(A)$.
@@
    This is an adaptation of Definition 2.1 on p. 575 of [18] by Hetyei which specializes to the original definition when $A$ is a simplicial complex. We can think of pairs of actual vertices of $A$ as vertices subdividing the corresponding edges and the pairs ending with $\widehat{1}$ correspond to vertices of the original cell complex $A$. The remaining conditions ensure that the “realization” lies inside some face of $A$ and we’re moving forward with respect to an edge ordering. Note that we think about faces of the subdivision in terms of which “side” of each edge we’re on. This is used later in the $f$-vector count of Tchebyshev subdivisions and is reminiscent of how one thinks about faces in barycentric subdivisions.
@@
    While each face $G$ of $T(A)$ of dimension $k-1$ has the “right” number of vertices to be a $(k-1)$-dimensional face of a simplicial complex, $T(A)$ itself isn’t necessarily a simplicial complex. This is because removing the underlying set of a vertex of $G$ (given by a singleton or a pair) from the face of $A$ given by the union of the underlying sets of the vertices of $G$ might not yield a face of $A$ if $A$ isn’t a simplicial complex.
@@
    For $k\geq 1$, the number of $(k-1)$-faces of of $T(A)$ is given by
@@
    $u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k}\}\setminus\{\widehat{1}\},$
    the rest of the work has to do what terms to pair together and in what order the pairs should be listed. The main structure of the proof has to do with the smallest face of the original cell complex $A$ that “contains” the given face of the Tchebyshev subdivision $T(A)$ and keeping track of possible slots for $u_{i}$’s and $v_{j}$’s (labeled with $+$ and -- signs), and how many strings with a given set of signs satisfy condition b) in Part 1 of Definition 2.26. This is sort of similar to how we keep track of faces on a barycentric subdivision.
@@
We will now consider the Tchebyshev triangulation (Definition 2.26) of the signed unused color complex from Definition 2.19. The following is a reproduction of Proposition 2.20 which will be relevant to our analysis.
Proposition 2.31.
Let $\Gamma$ be a balanced simplicial complex of dimension $D-1$. A collection of vertices $v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\in V(\mathcal{D}(\Gamma))$ with $\varepsilon_{i}=\pm 1$ forms a face of $\mathcal{D}(\Gamma)$ if and only if one of the following conditions (in the notation of Definition 2.19) holds:
  1. 1.
    $F_{1}\cap\cdots\cap F_{k}\neq\emptyset$.
  2. 2.
Note that each $C_{i}$ is determined by $F_{i}\in\Gamma$.
@@
$\frac{g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1=F(T(\mathcal{D}(\Gamma)),w)=f\left(T(\mathcal{D}(\Gamma)),\frac{w-1}{2}\right)=\frac{1}{2^{n(T(\mathcal{D}(\Gamma)))}}\widetilde{f}(M(T(\mathcal{D}(\Gamma))),w-1).$
@@
is $f_{k-1}(T(\mathcal{D}(\Gamma)))2^{-k}$. A $k$-tuple of vertices of $T(\mathcal{D}(\Gamma))$ forming a face of $\mathcal{D}(\Gamma)$ has between $k$ and $2k$ choices of signs to make for each vertex of the Tchebyshev subdivision from possible sizes in a cell complex $A$ of the union of vertices of $A$ coming from $k$ vertices of of the Tchebyshev subdivision $T(A)$ (analogous to considerations in the proof of Theorem 3.1 on p. 577 – 578 of [18]). This bound comes from having $2$ sign choices for every “valid” singleton of the form $(u,\widehat{1})$ from Definition 2.26 and $4$ sign choices when we have a valid pair $(u,v)$. This means that we have $2^{m}$ sign choices for some $k\leq m\leq 2k$. Effectively, what this means is that we’re considering $k$-tuples consisting of unsigned pairs of the form $(u,\widehat{1})$ and pairs $(u,v)$ where only the second coordinate is signed. Equivalently, $u$ and $v$ must have the same sign. This is a sort of “complex of one-sided intervals” $\widetilde{T}(\mathcal{D}(\Gamma))$ (a modification of the Tchebyshev subdivision $T(\mathcal{D}(\Gamma))$ of $\mathcal{D}(\Gamma)$ which we define below (with relevant parts reproduced from Definition 2.26).
@@
    $\{(u,v)\in V(\mathcal{D}(\Gamma)):u<v,\{u,v\}\subset F\text{ for some }F\in\mathcal{D}(\Gamma)\text{ and }\operatorname{sgn}(u)=\operatorname{sgn}(v)\}\cup\{(u,\widehat{1}):u\in V(\mathcal{D}(\Gamma))\}.$
  • The $(k-1)$-faces of $\widetilde{T}(\mathcal{D}(\Gamma))$ are all the sets $\{(u_{1},v_{1}),\ldots,(u_{k},v_{k})\}$ satisfying $v_{1}<\cdots<v_{k}$ and the following conditions:
@@
$[w^{k}]g_{\Delta}(2w)=\sum_{j=k}^{\frac{d}{2}}f_{j-1}(T^{*})2^{j}\binom{j}{k}$
@@
$[w^{k}]\frac{g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1=F(T(\mathcal{D}(\Gamma)),w)=\sum_{j=k}^{\frac{d}{2}}f_{j-1}(A)2^{j}\binom{j}{k},$
where $A$ denotes the $A(w)$ such that $R(w)=A(w+1)$. We can take $A(w)$ to have nonnegative coefficients when $\gamma(\Delta)=f(T)$.
@@
$[w^{k}]g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)-2\sum_{i=0}^{\frac{d}{2}-1}f_{\frac{d}{2}-i-1}(\Gamma)=2F(T(\mathcal{D}(\Gamma)),w)$ $=2\sum_{j=k}^{\frac{d}{2}}f_{j-1}(A)2^{j}\binom{j}{k}$
$\Longrightarrow[w^{k}]g_{\Delta}(2w)-\left(f_{\frac{d}{2}-1}(\Gamma)+2\sum_{i=1}^{\frac{d}{2}-1}f_{\frac{d}{2}-i-1}(\Gamma)\right)$ $=2\sum_{j=k}^{\frac{d}{2}}f_{j-1}(A)2^{j}\binom{j}{k},$
which means that we put $\pm$ signs on each face (alternatively put signs on an extraneous element $\widehat{1}$ for each face).
To obtain a cell complex $T^{*}$ where
$[w^{0}]g_{\Delta}(2w)=2\sum_{j=0}^{\frac{d}{2}}f_{j-1}(A)+\left(f_{\frac{d}{2}-1}(\Gamma)+2\sum_{i=1}^{\frac{d}{2}-1}f_{\frac{d}{2}-i-1}(\Gamma)\right)=\sum_{j=0}^{\frac{d}{2}}f_{j-1}(T^{*}),$
it suffices to formally “add” $\left(f_{0}(\Gamma^{*})+2\sum_{i=1}^{\frac{d}{2}-1}f_{i}(\Gamma^{*})\right)$ disjoint points since $f_{0}(T^{*})$ does not appear in the sums determining $[w^{k}]g_{\Delta}(2w)$ for any $k\geq 1$.
@@
Under the modifications above, the mirroring/Danzer complex $M(T(\mathcal{D}(\Gamma)))$ is the (signed) interval complex/generalized cubical barycentric subdivision associated to the (dual of) a simplicial complex $S$ such that $\gamma(\Delta)=f(S)$ if one exists.
Finally, we study some geometric properties connected to nonnegativity of translation by $-1$ (thinking about the gamma vector). Given $MA$, note that the links of the vertices are isomorphic to $A$. The following result shows that $MA$ is a $CAT(0)$ cubical complex if and only if $A$ is a flag simplicial complex and $MA$ is simply connected:
@@
Every flag simplicial complex is the crossing complex $\Delta_{P}$ (associated to a $PIP$ $P$) of some $CAT(0)$ cubical complex.
@@
The map $P\mapsto\Box_{P}$ is a bijection from the set of $PIP$s to the set of rooted $CAT(0)$ cubical complexes, with inverse given by $\Box\mapsto P_{\Box}$.

### v2

@@ 1.2.2 Colorings, Tchebyshev subdivisions, and intervals
$[\alpha^{k}]\frac{P_{\Delta}(2\alpha+1)+h_{\frac{d}{2}}(\Delta)}{2}$ $=[\alpha^{k}]\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}$
$=\sum_{j=k}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)2^{k}\binom{j}{k}$
$=2^{k}\sum_{j=k}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)\binom{j}{k}$
for $k\geq 1$ and the constant term of
$\frac{P_{\Delta}(2\alpha+1)+h_{\frac{d}{2}}(\Delta)}{2}=\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}$
is
$[\alpha^{0}]\frac{P_{\Delta}(2\alpha+1)+h_{\frac{d}{2}}(\Delta)}{2}$ $=[\alpha^{0}]\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}$
$=\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma).$
Given $k\geq 1$, we claim that the $f$-vector of the complex $\mathcal{D}(\Gamma)$ from Definition 1.22 has dimensions given by the sum above. Given a $\left(\frac{d}{2}-j-1\right)$-dimensional face $F\in\Gamma$, there are $j$ unused colors forming the set $C_{F}\subset\left[\frac{d}{2}\right]$. Choosing $k$ of them gives the subset $Q\subset C_{F}$ from Definition 1.22 and the number of pairs $(F,Q)\in\Gamma\times\left[\frac{d}{2}\right]$ with $\dim F=\frac{d}{2}-j-1$ and $Q\subset C_{F}$ is counted by the term $f_{\frac{d}{2}-j-1}(\Gamma)\binom{j}{k}$ in the sum. Assigning a $\pm$ signs to each element of $Q$ gives the multiplication by $2^{k}$.
@@
(Definition 3.2 on p. 578) Given a cell complex (Definition 1.3) $A$ of dimension D1D-1, the $F$-polynomial of $A$ is
@@
Given a cell complex $A$ such that (k1)(k-1)-dimensional faces have $k$ vertices, we have that
@@
for a cell complex $T$ with $n$ vertices such that (k1)(k-1)-dimensional faces have $k$ vertices.
@@
(Cell complex generalization of Definition 2.1 on p. 573 of [19])
Let $A$ be (the dual of) a cell complex from Definition 1.3.
Given a cell complex $A$ on a vertex set $V$, fix a linear order << on $V$ and introduce a new element $\widehat{1}$ larger than all elements of $V$. The Tchebyshev triangulation $T_{<}A$ of $A$ is the cell complex on the vertex set
@@
whose (k1)(k-1)-faces are all sets $\{(u_{1},v_{1}),\ldots,(u_{k},v_{k})\}$ satisfying $v_{1}<\cdots<v_{k}$ and the following conditions:
@@
Since the $f$-vector of $T_{<}(A)$ is independent of the ordering <<, we will often drop the ordering and simply write $T(A)$.
@@
    This is an adaptation of Definition 2.1 on p. 573 of [19] by Hetyei which specializes to the original definition when $A$ is a simplicial complex. We can think of pairs of actual vertices of $A$ as vertices subdividing the corresponding edges and the pairs ending with $\widehat{1}$ correspond to vertices of the original cell complex $A$. The remaining conditions ensure that the “realization” lies inside some face of $A$ and we’re moving forward with respect to an edge ordering. Note that we think about faces of the subdivision in terms of which “side” of each edge we’re on. This is used later in the $f$-vector count of Tchebyshev subdivisions and is reminiscent of how one thinks about faces in barycentric subdivisions.
@@
    While each face $G$ of $T(A)$ of dimension k1k-1 has the “right” number of vertices to be a (k1)(k-1)-dimensional face of a simplicial complex, $T(A)$ itself isn’t necessarily a simplicial complex. This is because removing the underlying set of a vertex of $G$ (given by a singleton or a pair) from the face of $A$ given by the union of the underlying sets of the vertices of $G$ might not yield a face of $A$ if $A$ isn’t a simplicial complex.
@@
    For $k\geq 1$, the number of (k1)(k-1)-faces of of $T(A)$ is given by
@@
    $\{u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k}\}\setminus\{\widehat{1}\}$,
    the rest of the work has to do what terms to pair together and in what order the pairs should be listed. The main structure of the proof has to do with the smallest face of the original cell complex $A$ that “contains” the given face of the Tchebyshev subdivision $T(A)$ and keeping track of possible slots for $u_{i}$’s and $v_{j}$’s (labeled with ++ and - signs), and how many strings with a given set of signs satisfy condition b) in Part 1 of Definition 1.29. This is sort of similar to how we keep track of faces on a barycentric subdivision.
@@
We will now consider the Tchebyshev triangulation (Definition 1.29) of the signed unused color complex from Definition 1.22.
@@
$\frac{g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1$ $=f\left(T(\mathcal{D}(\Gamma)),\frac{w-1}{2}\right)$
$=\frac{1}{2^{n(T(\mathcal{D}(\Gamma)))}}\widetilde{f}(M(T(\mathcal{D}(\Gamma))),w-1).$
@@
[…]
@@
    $\{(u,v)\in V(\mathcal{D}(\Gamma)):u<v,\{u,v\}\subset F\text{ for some }F\in\mathcal{D}(\Gamma)$
    and
    $\operatorname{sgn}(u)=\operatorname{sgn}(v)\}\cup\{(u,\widehat{1}):u\in V(\mathcal{D}(\Gamma))\}.$
  • The (k1)(k-1)-faces of $\widetilde{T}(\mathcal{D}(\Gamma))$ are all the sets $\{(u_{1},v_{1}),\ldots,(u_{k},v_{k})\}$ satisfying $v_{1}<\cdots<v_{k}$ and the following conditions:
@@
$w^{k}]g_{\Delta}(2w)=\sum_{j=k}^{\frac{d}{2}}f_{j-1}(T^{*})2^{j}\binom{j}{k}$
@@
$[w^{k}]\left(\frac{g_{\Delta}(2w)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1\right)$ $=[w^{k}]F(T(\mathcal{D}(\Gamma)),w)$
$=[w^{k}]R(w-1)$
$=[w^{k}]A(w),$
where $A$ denotes the polynomial $A(w)$ such that $R(w)=A(w+1)$ ($R(w)$ itself defined below Corollary 1.33). We can take $A(w)$ to have nonnegative coefficients if $\gamma(\Delta)=f(T)$.
If we substitute $w=\mu+1$, we have
$[\mu^{k}]\left(\frac{g_{\Delta}(2(\mu+1))+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1\right)$ $=[\mu^{k}]F(T(\mathcal{D}(\Gamma)),\mu+1)$
$=[\mu^{k}]R(\mu)$
$=[\mu^{k}]A(\mu+1)$
$=\sum_{j=k}^{\frac{d}{2}}A_{j-1}\binom{j}{k}$,
where $A_{j-1}$ denotes the coefficient of $x^{j}$ in $A(x)$. Note that
$\mu^{k}]R(\mu)=[\mu^{k}]f\left(T(\mathcal{D}(\Gamma)),\frac{\mu}{2}\right)=f_{k-1}(T(\mathcal{D}(\Gamma)))2^{-k}=f_{k-1}(\widetilde{T}(\mathcal{D}(\Gamma)))$
and
$g_{\Delta}(2w)$ $=(2w+2)^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2w+2}\right)$
$\Longrightarrow g_{\Delta}(2(\mu+1))$ $=(2(\mu+1)+2)^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2(\mu+1)+2}\right)$
$=(2(\mu+2))^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2(\mu+2)}\right)$
$=(2\mu+4)^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2\mu+4}\right).$
@@
$[\mu^{k}]\left(g_{\Delta}(2(\mu+1))+f_{\frac{d}{2}-1}(\Gamma)-2\sum_{i=0}^{\frac{d}{2}}f_{\frac{d}{2}-i-1}(\Gamma)+1\right)$ $=2[\mu^{k}]F(T(\mathcal{D}(\Gamma)),\mu+1)$
$=2[\mu^{k}]R(\mu)$
$=2\cdot 2^{k}[\mu^{k}]f(\widetilde{T}(\mathcal{D}(\Gamma)),\mu)$
$=2\cdot 2^{k}[\mu^{k}]\widetilde{A}(\mu+1)$
$=2\cdot 2^{k}\sum_{j=k}^{\frac{d}{2}}\widetilde{A}_{j-1}\binom{j}{k}$,
where $\widetilde{A}(x)=f(\widetilde{T}(\mathcal{D}(\Gamma)),x-1)$ and $\widetilde{A}_{j-1}$ denotes the coefficient of $x^{j}$ in $\widetilde{A}(x)$. The multiplication by $2^{k}$ means that we put $\pm$ signs on each face (alternatively put signs on an extraneous element $\widehat{1}$ for each face).
Suppose that $\widetilde{A}(x)$ has nonnegative coefficients. To obtain a cell complex $T^{*}$ where
$w^{0}]g_{\Delta}(2w)=2\sum_{j=0}^{\frac{d}{2}}f_{j-1}(A)+\left(f_{\frac{d}{2}-1}(\Gamma)+2\sum_{i=1}^{\frac{d}{2}-1}f_{\frac{d}{2}-i-1}(\Gamma)\right)=\sum_{j=0}^{\frac{d}{2}}f_{j-1}(T^{*}),$
it suffices to formally “add” $\left(f_{0}(\Gamma^{*})+2\sum_{i=1}^{\frac{d}{2}-1}f_{i}(\Gamma^{*})\right)$ disjoint points since $f_{0}(T^{*})$ does not appear in the sums determining $w^{k}]g_{\Delta}(2w)$ for any $k\geq 1$.
@@
  1. 1.
    Suppose that $h(\Delta)$ is the $h$-vector of a flag sphere of odd dimension d1d-1 (i.e. $d$ is even). If $d\geq 4$, then the polynomial $g_{\Delta}(2(u+1))=(2(u+2))^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2(u+2)}\right)$ has nonnegative coefficients. After multiplying by a power of 2, we have the $f$-polynomial of the mirroring/Danzer/power complex $M(T(\mathcal{D}(\Gamma)))$ with a nonnegative constant added to it.
  2. 2.
    By the $f_{k}^{\Box}(S)$ formula below Definition 1.12, substituting the variable $w$ by $u+1$ sends the $f$-polynomial of a simplicial complex to that of its cubical barycentric subdivision. Consider the $f$-polynomial of the mirroring/Danzer/power complex $M(T(\mathcal{D}(\Gamma)))$.
    Corollary 1.33 implies that its $f$-polynomial behaves “formally” like a double cubical barycentric subdivision associated to some integer tuple. If $\gamma(\Delta)=f(T)$, we can remove the word “formal” and take the $f$-polynomial associated to a cell complex after a rescaling as the input.
Finally, we study some geometric properties connected to nonnegativity of translation by 1-1 (thinking about the gamma vector). Given $MA$, note that the links of the vertices are isomorphic to $A$. The following result shows that $MA$ is a $CAT(0)$ cubical complex if and only if $A$ is a flag simplicial complex and $MA$ is simply connected:
@@
Every flag simplicial complex is the crossing complex $\Delta_{P}$ (associated to a PIPPIP $P$) of some $CAT(0)$ cubical complex.
@@
The map $P\mapsto\Box_{P}$ is a bijection from the set of PIPPIPs to the set of rooted $CAT(0)$ cubical complexes, with inverse given by $\Box\mapsto P_{\Box}$.